The electron and proton of a hydrogen atom are separated (on the average) by a distance of approximately 5.3x10^-11 m. Find the magnitude of the electric force and the gravitational force between the two particles


1
Expert's answer
2022-02-13T12:15:35-0500

Given:

"r=5.3*10^{-11}\:\rm m"

"e=1.6*10^{-19}\:\rm C"

"m_e=9.1*10^{-31}\:\rm kg"

"m_p=1.67*10^{-27}\;\rm kg"


The magnitude of the electric force

"F_e=k\frac{q_1q_2}{r^2}=k\frac{e^2}{r^2}\\
=\rm 9*10^9*\frac{(1.6*10^{-19})^2}{(5.3*10^{-11})^2}=8.2*10^{-8}\:N"

The magnitude of the gravitational force

"F_g=G\frac{m_1m_2}{r^2}\\
=\rm 6.67*10^{-11}*\frac{1.67*10^{-27}*9.1*10^{-31}}{(5.3*10^{-11})^2}=3.6*10^{-47}\:N"